OpenStudy (anonymous):

Continuous random variable X has density given by f(x)=c⋅(x+1)*(x−3) if x∈[−1,3] and 0 otherwise. 1) Find DF(2;X) . distribution function 2) Find E[X] . C= -3/32

OpenStudy (amistre64):

what does 2;x mean?

OpenStudy (anonymous):

is something about the notation of x>2 and x<=2

OpenStudy (amistre64):

i see notifs arent posting ...

OpenStudy (anonymous):

i guess i need to put hat in the integral

OpenStudy (amistre64):

im not used to the notation used, can you provide a link to what explains it? otherwise i wouldnt know if my input is worth anything.

OpenStudy (anonymous):

is not in the site rather is on my notes

OpenStudy (amistre64):

and i would assume since your distribution is a quadratic, that its symmetry would give us the vertex as its mean

OpenStudy (anonymous):

perpaps

OpenStudy (amistre64):

density = P(x) distribution = \(\int P(x)\)

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

i think is support to be like this ? \[\int\limits_{-2}^{2}\]

OpenStudy (amistre64):

lower limit doesnt go past -1, everything after that is 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i guess there is no way how to solved this :(

OpenStudy (anonymous):

here are the notes :(

OpenStudy (anonymous):

i will close this one open a new one this is kind of a mess :(